OpenStudy (anonymous):

which fraction is NOT equivalent to 4\12? A. 1\3 B. 12\30 C. 8\24 D. 16\48

OpenStudy (turingtest):

B 4/12=1/3 12/30=2/5

OpenStudy (anonymous):

simplify each of them including 4/12 4/12=1/3 so see which one does not equal 1/3 or 4/12

OpenStudy (anonymous):

thanks sooo much :)

OpenStudy (turingtest):

you're welcome but medals are better